OpenStudy (anonymous):

Which amusement park ride utilizes periodic motion? A. slide B. Bumper cars C. giant swing D. roller coaster

OpenStudy (anonymous):

C, as it's a pendulum, they oscillate, have a time period.

OpenStudy (anonymous):

\[T \approx 2\pi \sqrt{\frac{ L }{ g }}\]